• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Z. E

    Log entry time 00:01:39 on December 9,2008

    Entry number 252602

    keyword=Shift Summary

    16:00 Shift start
    Moller measurement in progress
    19:30 Calling MCC to restore the standard tune for production.
    20:10 after tune on empty put BeO target in
    20:17 MCC called they need more time for tuning.
    20:30 Beam on Beo at 3uA positions on BPMs good and Spot++ check good
    20:50 Moved the target to He3 at 2uA raster 3X2.5 MCC(3X4), spot++ good
    21:00 Raster MCC(5x6.6) spot++ looks good
    21:10 Raster back to MCC (3x4) check spot++ again
    21:15 Bigbite HVs and RICH detector HVs on

    21:18 Production runs start with run 4305
    21:20-21:51 run # 4305 0.5M evts, good, charge: 0.02139
    21:40 Hall B measure also a low polarization of the beam consistent with Eugene's Moller measurement.
    21:50 Beam stopped the Electron Gun Group (EGG) is in to check the polarization of the beam. Pockel cells voltage is suspected.
    22:59 MCC starts sending tune beam to our hall again
    They think the problem is resolved. They fixed the voltages on the pockel cells. Of course we need to take another polarization measurement.
    23:30 Checked beam raster at 2uA on He target.
    BPMA (0.10,1.7) BPMB (-1.5,1.2)
    Start production run # 4308
    23:37-.... run 4308
    24:00 Shift Ends during run 4308